Given the following data:

Fe2O3(s) + 3 CO(g) → 2 Fe(s) + 3 CO2(g)      ΔH = -26.8 kJ

   FeO(s) + CO(g) → Fe(s) + CO2(g)             ΔH = -16.5 kJ

Calculate the ΔH for the reaction below:

Fe2O3(s) + CO(g) → 2 FeO(s) + CO2(g)


1
Expert's answer
2021-02-04T07:07:41-0500

Solution:

Given the chemical reaction:

Fe2O3(s) + 3 CO(g) → 2 Fe(s) + 3 CO2(g)      ΔH = -26.8 kJ

FeO(s) + CO(g) → Fe(s) + CO2(g)             ΔH = -16.5 kJ


Firstly we should find the standard enthalpies of Fe2O3(s) and FeO(s) in terms of Hess`s law:

"\Delta H_{reaction}=\Delta H_{product}-\Delta H_{reactant}"


"\Delta H(Fe_2O_3) = -822.2 (\frac{kJ}{mole})"


"\Delta H(FeO)=-264.8 (\frac{kJ}{mole})"


Then, we put these values to the main chemical reaction:


"Fe_2O_3+CO\rightarrow2FeO+CO_2"


Hess`s law:


"\Delta H_{reaction}=\Delta H_{product}-\Delta H_{reactant}"


"\Delta H_{product}=" 2(-264.8 "(\frac{kJ}{mole})" ) + (-393.5 "(\frac{kJ}{mole})" )= -923.1 kJ


"\Delta H_{reactant}=" (-822.2 "(\frac{kJ}{mole})" ) + (-110.5 "(\frac{kJ}{mole})" ) = -932.7 kJ


"\Delta H_{reaction}=" (-923.1 kJ) - ( -932.7 kJ) = 9.6 kJ


Answer:

"\Delta H_{reaction}=9.6kJ"